A chemist measures the energy change Δ H during the following reaction: 2 NO 2 (g) → N 2 O 4 (g) = Δ H − 55.3 kJ Use the information to answer the following questions.

his reaction is...

endothermic.

exothermic.

Suppose 97.0g of NO2 react. Will any heat be released or absorbed?

Yes, absorbed.

Yes, released.

No.

If you said heat will be released or absorbed in the second part of this question, calculate how much heat will be released or absorbed.

kJ

Be sure your answer has the correct number of significant digits.
Answer #1

1)
since ΔH is negative, the reaction is exothermic
Answer: exothermic

2)
Reaction is exothermic
SO, heat is being released
Answer:
yes, released

3)

Molar mass of NO2,
MM = 1*MM(N) + 2*MM(O)
= 1*14.01 + 2*16.0
= 46.01 g/mol


mass(NO2)= 97.0 g

use:
number of mol of NO2,
n = mass of NO2/molar mass of NO2
=(97 g)/(46.01 g/mol)
= 2.108 mol
Since Δ H is negative, heat is released
when 2 mol of NO2 reacts, heat released = 55.3 KJ
So,
for 2.108 mol of NO2, heat released = 2.108*55.3/2 KJ
= 58.29 KJ
Answer: 58.3 KJ